OpenStudy (anonymous):

√80 Simplify each radical expression. -- √10

OpenStudy (john_es):

\[\frac{\sqrt{80}}{\sqrt{10}}=\frac{\sqrt{8}}{\sqrt{1}}=2\sqrt{2}\]